# PerCoR: Evaluating Commonsense Reasoning in Persian via Multiple-Choice Sentence Completion

### v1

@@ Abstract
We introduced PerCoRPersian Commonsense Reasoning—the first large-scale Persian benchmark for commonsense reasoning. PerCoR contains 106K multiple-choice sentence-completion problems drawn from more than forty news, cultural and other web sources. We introduce a novel conjunction-based segmentation strategy to generate coherent sentence–completion pairs, enabling broad topical and structural diversity. To create challenging distractors, we propose DRESS-AFDistractor Ranking via Embedding Similarity Scoring and Adversarial Filtering—a generation-free adversarial filtering method that selects distractors from the pool of gold continuations while maximising model confusion. Human annotators score 89% on PerCoR, while OpenAI-o3 achieves the highest performance at 92.18%, followed closely by Claude-Sonnet-3.7 (91.17%). The strongest open-source model, DeepSeek-R1, reaches 82.51%, underscoring both the dataset’s difficulty and the remaining performance gap in Persian commonsense reasoning. We further show that DRESS-AF transfers to the English HellaSwag benchmark, increasing its difficulty without hurting human solvability. The dataset is available at https://huggingface.co/datasets/MCINext/PerCoR.
@@
Morteza Alikhani   Mohammadtaha Bagherifard†∗   Erfan Zinvandi   Mehran Sarmadi MCINEXT {morteza.alikhani95, e.zeynvandi1376, mehran.sarmadi99}@sharif.edu taha.bagheri98@gmail.com
@@ 1 Introduction
Despite recent progress in Persian NLP through resources such as ParsiNLU (Khashabi et al., 2021), PersianQA (Ayoubi, 2021), and Pquad (Darvishi et al., 2023), Persian remains a low-resource language for high-level reasoning tasks, particularly commonsense inference. This leaves a significant gap in evaluating and advancing structured reasoning capabilities in this language. To address this limitation, we introduce PerCoR—the first large-scale Persian commonsense reasoning dataset in multiple-choice sentence–completion format. Constructed from over 40 diverse Persian websites, PerCoR captures a broad range of domains and linguistic styles. We formulate each instance as a sentence prefix followed by four completion candidates: one correct and three distractors. Instead of relying on simple rule-based methods for sentence segmentation, we generate sentence–completion pairs by splitting at conjunctions, promoting natural flow and semantic coherence. Unlike oversimplified strategies such as the one employed in SWAG (Zellers et al., 2018), which relies on temporally grounded data like video captions, our conjunction-based approach is applicable to a wide range of textual sources. This enables broader domain coverage and greater variability in sample length, enhancing both the diversity and richness of the dataset.
@@ 2 Related Work
Numerous English benchmarks have been introduced to evaluate commonsense reasoning in multiple formats. SWAG (Zellers et al., 2018) and HellaSwag (Zellers et al., 2019) pose multiple-choice sentence completion tasks based on narrative or descriptive contexts. HellaSwag, in particular, uses adversarial filtering to create distractors that are challenging for language models but easily solvable by humans. Other benchmarks such as WinoGrande (Sakaguchi et al., 2021), CommonsenseQA (Talmor et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), Cosmos (Huang et al., 2019), and Social IQa (Sap et al., 2019) cover a variety of commonsense dimensions, including physical reasoning, social dynamics, and multi-hop inference. More recent efforts include GLUCOSE (Mostafazadeh et al., 2020), a dataset of causal explanations in short narratives, annotated across ten dimensions of inferential knowledge; COM2SENSE (Singh et al., 2021), which evaluates a model’s ability to discriminate between true and false commonsense statements in complementary pairs; and CommonsenseQA 2.0 (Talmor et al., 2021), an adversarially curated yes/no question dataset designed to be difficult for large language models while remaining easy for humans. Despite substantial progress, these benchmarks are primarily designed for English, leaving a gap in resources for other languages.
@@ 4.2.2 HellaSwag Dataset
While increasing dataset difficulty is desirable, it is crucial that the dataset remains answerable and reliable. To assess this, we conduct a human evaluation on a 200-sample subset from each dataset. Each dataset is annotated by a human annotator. The resulting human accuracies are 90% for the original HellaSwag, 89.5% for the easier variant, and 83% for the harder variant. Although the trend mirrors the degradation observed in model performance, the drop in human accuracy is modest by comparison. These results highlight an important point: unlike LLM-generated distractors that may introduce stylistic or fluency biases, our embedding-based distractor construction avoids generation artifacts, allowing humans to perform more consistently. This suggests that while DRESS-AF introduces additional challenge, it maintains dataset integrity. In sum, DRESS-AF yields more difficult yet human-solvable distractors, effectively benchmarking model robustness without compromising dataset quality.
@@ 4.3 Model Results on PerCoR
Overall, closed‐source models dominate: OpenAI-o3 (OpenAI, 2025b) tops the leaderboard at 92.18 %, followed by Claude-3.7-Sonnet (Anthropic, 2025) (91.17 %) and GPT‑4.1 (OpenAI, 2025a) (88.39 %); the best open‐source checkpoint, DeepSeek‑R1 (DeepSeek-AI et al., 2025a), reaches 82.51 %, trimming the gap to roughly 10%, while most open‐source peers fall between 60 % and 80 %. Human majority‐vote accuracy on PerCoR is 89 % (details in Appendix B.4), so only o3 and Sonnet currently exceed non‐expert annotators performance. Despite strong aggregate performance, top-performing models still exhibit occasional failures on nuanced reasoning cases—several examples are provided in Appendix B.6.
@@
In summary: (i) PerCoR is a challenging benchmark—only two proprietary models exceed 90 % accuracy, while the best open-source model, DeepSeek-R1, still lags by $\sim$10 %; (ii) post-processing plays a crucial role in revealing latent reasoning capabilities, especially for models that embed correct answers in natural language rather than the required format; (iii) reasoning-oriented fine-tuning and alignment are key—OpenAI-o3 leads all models, and DeepSeek-R1 outperforms DeepSeek-V3 in strict accuracy, highlighting its superior adherence to the expected output format; and (iv) there remains ample headroom for open-source models to close the gap—not only through better prompt-following and format alignment, but also via lightweight, resource-efficient fine-tuning; our adaptation of these models with limited data surpassed the strongest open-source zero-shot baselines, highlighting the impact of minimal task-specific supervision.
@@ 5 Conclusion
We introduced PerCoR, a 106K-example benchmark that fills a major evaluation gap for commonsense reasoning in Persian. Our conjunction-based extraction strategy generates natural sentence-completion pairs from static prose, while DRESS-AF produces hard, language-agnostic distractors without resorting to LLM generation. Benchmarking 32 models reveals a persistent ten-point gap between the strongest open and closed systems, and qualitative analysis highlights residual weaknesses in discourse-level reasoning.
@@ B.4 Human Evaluation
We used Label Studio Tkachenko et al. (2020-2022) to evaluate the accuracy on the test split of the dataset. Each sample was annotated independently by three human annotators. In cases where at least two annotators agreed on the same label, their consensus was taken as the final label, which was then compared with the provided label. If all three annotators disagreed, the sample was considered incorrectly labelled. Importantly, annotators worked independently and were not aware of each other’s selections.
@@ B.5 Model Fine-tuning on the Dataset
B.6 Qualitative Failure Cases
Despite these strong overall results, even top-performing closed-source models occasionally fail on examples requiring subtle syntactic, temporal, or discourse-level reasoning. Several illustrative failure cases are provided in Figures 1417, where the selected completions are semantically or grammatically incoherent. Each example is accompanied by an explanation clarifying why the chosen option is incorrect and why the correct answer better satisfies the continuation constraints. These qualitative insights highlight that high aggregate accuracy can mask nuanced reasoning failures that warrant deeper analysis.

### v2

@@ Abstract
We introduced PerCoRPersian Commonsense Reasoning—the first large-scale Persian benchmark for commonsense reasoning. PerCoR contains 106K multiple-choice sentence-completion problems drawn from more than forty news, cultural and other web sources. We adopt a linguistically grounded, conjunction-based segmentation strategy to generate coherent prefix–continuation pairs. To create challenging distractors, we propose DRESS-AFDistractor Ranking via Embedding Similarity Scoring and Adversarial Filtering—a generation-free adversarial filtering method that selects distractors from the pool of gold continuations while maximising model confusion. Human annotators score 89% on PerCoR, while OpenAI-o3 achieves the highest performance at 92.18%, followed closely by Claude-Sonnet-3.7 (91.17%). The strongest open-source model, DeepSeek-R1, reaches 82.51%, underscoring both the dataset’s difficulty and the remaining performance gap in Persian commonsense reasoning. We further show that DRESS-AF transfers to the English HellaSwag benchmark, increasing its difficulty without hurting human solvability. The dataset is available at https://huggingface.co/datasets/MCINext/PerCoR.
@@
Morteza Alikhani   Mohammadtaha Bagherifard†∗   Erfan Zinvandi Mehran Sarmadi MCINEXT taha.bagheri98@gmail.com {morteza.alikhani95, e.zeynvandi1376, mehran.sarmadi99}@sharif.edu
@@ 1 Introduction
Despite recent progress in Persian NLP through resources such as ParsiNLU (Khashabi et al., 2021), PersianQA (Ayoubi, 2021), and Pquad (Darvishi et al., 2023), Persian remains a low-resource language for high-level reasoning tasks, particularly commonsense inference. This leaves a significant gap in evaluating and advancing structured reasoning capabilities in this language. To address this limitation, we introduce PerCoR—the first large-scale Persian commonsense reasoning dataset in multiple-choice sentence–completion format. Constructed from over 40 diverse Persian websites, PerCoR captures a broad range of domains and linguistic styles. We formulate each instance as a sentence prefix followed by four completion candidates: one correct and three distractors. Instead of relying on simple rule-based methods for sentence segmentation, we generate sentence–completion pairs by splitting at conjunctions, promoting natural flow and semantic coherence. This connective-driven formulation is related to prior work that uses discourse markers for inference and sentence completion (Bhargava and Ng, 2022), but PerCoR applies this strategy at scale to Persian and across diverse, non-narrative domains. Unlike oversimplified strategies such as the one employed in SWAG (Zellers et al., 2018), which relies on temporally grounded data like video captions, our conjunction-based approach is applicable to a wide range of textual sources. This enables broader domain coverage and greater variability in sample length, enhancing both the diversity and richness of the dataset.
@@
PerCoR specifically targets discourse-level and contextual commonsense reasoning. To select the correct continuation, a model must infer the semantic relation introduced by the conjunction (e.g., causal, contrastive, elaborative) and identify which candidate maintains coherence with the preceding context. Distractors are intentionally plausible in isolation but violate these discourse constraints, requiring models to integrate implicit world knowledge with local coherence cues rather than relying on shallow lexical matching.
@@ 2 Related Work
Numerous English benchmarks have been introduced to evaluate commonsense reasoning in multiple formats. SWAG (Zellers et al., 2018) and HellaSwag (Zellers et al., 2019) pose multiple-choice sentence completion tasks based on narrative or descriptive contexts. HellaSwag, in particular, uses adversarial filtering to create distractors that are challenging for language models but easily solvable by humans. Cloze-style narrative continuation tasks such as the Story Cloze Test (Mostafazadeh et al., 2016) also involve selecting a plausible continuation, though they focus on narrative coherence rather than connective-driven discourse inference. Other benchmarks such as WinoGrande (Sakaguchi et al., 2021), CommonsenseQA (Talmor et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), Cosmos (Huang et al., 2019), and Social IQa (Sap et al., 2019) cover a variety of commonsense dimensions, including physical reasoning, social dynamics, and multi-hop inference. More recent efforts include GLUCOSE (Mostafazadeh et al., 2020), a dataset of causal explanations in short narratives, annotated across ten dimensions of inferential knowledge; COM2SENSE (Singh et al., 2021), which evaluates a model’s ability to discriminate between true and false commonsense statements in complementary pairs; and CommonsenseQA 2.0 (Talmor et al., 2021), an adversarially curated yes/no question dataset designed to be difficult for large language models while remaining easy for humans. Despite substantial progress, these benchmarks are primarily designed for English, leaving a gap in resources for other languages.
@@ 4.2.2 HellaSwag Dataset
While increasing distractor difficulty is desirable, it is equally important that the dataset remain answerable and free of generation-induced artifacts. To assess this, we conduct a human evaluation on a 200-sample subset of each variant. On the original HellaSwag, human accuracy is 90%. This remains nearly unchanged on the easier variant (89.5%) and decreases only modestly on the harder one (83%). In contrast, model accuracy drops much more sharply: for Gemma-3-27B-it and GPT-4o-mini, performance decreases by 5–7 points on the easier set and by 14–16 points on the harder set (Figure 5). This asymmetry indicates that models were partially exploiting stylistic or fluency artifacts present in the original LM-generated distractors—artifacts that humans did not rely on. By selecting distractors from real human-written continuations, DRESS-AF reduces such generation-related biases and forces models to rely more on contextual and commonsense reasoning. At the same time, the modest change in human accuracy confirms that the resulting distractors remain human-solvable. In sum, DRESS-AF increases difficulty while maintaining dataset integrity and provides a more robust evaluation signal than generation-based AF methods.
@@ 4.3 Model Results on PerCoR
Overall, closed‐source models dominate: OpenAI-o3 (OpenAI, 2025b) tops the leaderboard at 92.18 %, followed by Claude-3.7-Sonnet (Anthropic, 2025) (91.17 %) and GPT‑4.1 (OpenAI, 2025a) (88.39 %); the best open‐source checkpoint, DeepSeek‑R1 (DeepSeek-AI et al., 2025a), reaches 82.51 %, trimming the gap to roughly 10%, while most open‐source peers fall between 60 % and 80 %. Human majority‐vote accuracy on PerCoR is 89 % (details in Appendix B.4), so only o3 and Sonnet currently exceed non‐expert annotators performance. Despite strong aggregate performance, top-performing models still exhibit occasional failures on nuanced reasoning cases—several examples are provided in Appendix.
@@
4.4 Qualitative Error Analysis
Despite strong overall performance, even top models occasionally fail on examples requiring subtle syntactic, temporal, or discourse-level reasoning. Appendix Figures 1417 present four illustrative failure cases in which the model-selected continuation is semantically implausible, grammatically incompatible, or violates the discourse relation implied by the conjunction. Each example includes a brief explanation of why the chosen option is incorrect and why the gold continuation better satisfies the coherence constraints. These qualitative observations reveal failure patterns that are not apparent from aggregate accuracy alone and highlight the value of PerCoR for probing nuanced reasoning and alignment behaviours in large language models.
4.5 Transferability to Persian QA
Prior work such as Social IQA (Sap et al., 2019) and WinoGrande (Sakaguchi et al., 2021) has demonstrated transferability by fine-tuning models on a commonsense benchmark and evaluating on other reasoning tasks (e.g., COPA (Gordon et al., 2011) and KnowRef (Emami et al., 2019)). Such evaluation is not currently feasible in Persian, as PerCoR is the first structured commonsense reasoning benchmark in the language. We hope PerCoR will serve as a foundation that enables similar cross-task studies in the future. To provide an initial signal of transferability under current resource constraints, we evaluate on PQUAD (Darvishi et al., 2023), a Persian extractive QA benchmark in which many queries require implicit inference and bridging context. We fine-tune Qwen3-32B using LoRA with a causal language modelling objective on PerCoR, then zero-shot evaluate both the base and fine-tuned models on the PQUAD test set (964 examples) using Exact Match (EM), ROUGE-L (Lin, 2004), and BLEU (Papineni et al., 2002).
Fine-tuning on PerCoR yields a small improvement in EM (25.21→26.56), but substantially increases ROUGE-L (19.27→29.69, +10.4) and BLEU (16.00→23.98, +7.9). These gains suggest that the discourse-level coherence learned from PerCoR transfers to extractive QA settings where models must integrate contextual cues with implicit knowledge. We view this as an encouraging early signal of transferability and include it to underscore PerCoR’s utility beyond its primary role as an evaluation benchmark.
In summary: (i) while PerCoR is approachable for top proprietary systems, it remains challenging for open-source models—only two models exceed 90 % accuracy, whereas the strongest open-source model still trails by approximately ten points; (ii) post-processing is often essential for revealing latent reasoning ability in models that express answers in free-form text rather than the required label format; (iii) reasoning-oriented alignment matters—OpenAI-o3 leads all systems, and DeepSeek-R1 surpasses DeepSeek-V3 in strict accuracy due to stronger format adherence; (iv) lightweight task-specific fine-tuning remains highly effective, with small adaptations outperforming the strongest open-source zero-shot baselines; (v) even top-performing models exhibit systematic mistakes on examples requiring subtle syntactic, temporal, or discourse-level reasoning, underscoring PerCoR’s diagnostic value; and (vi) the discourse-level coherence captured by PerCoR shows early signs of transferability to other Persian tasks, as reflected in improvements on PQUAD.
@@ 5 Conclusion
We introduced PerCoR, a 106K-example benchmark that fills a major evaluation gap for commonsense reasoning in Persian. Our conjunction-based extraction strategy yields natural sentence–completion pairs from diverse prose, and DRESS-AF selects challenging, human-written distractors while avoiding stylistic artifacts associated with LLM-generated options. Evaluating 32 models reveals a persistent ten-point gap between the strongest open-source and proprietary systems, while qualitative analysis exposes systematic failures on examples requiring subtle syntactic, temporal, or discourse-level reasoning. Together with early signs of transferability to downstream Persian tasks, these results show that PerCoR provides a robust benchmark, a fine-grained diagnostic tool, and a valuable resource for transfer learning in low-resource settings.
@@ B.4 Human Evaluation
We used Label Studio Tkachenko et al. (2020) to evaluate the accuracy on the test split of the dataset. Each sample was annotated independently by three human annotators. In cases where at least two annotators agreed on the same label, their consensus was taken as the final label, which was then compared with the provided label. If all three annotators disagreed, the sample was considered incorrectly labelled. Importantly, annotators worked independently and were not aware of each other’s selections.